# Bayesian Prediction under Moment Conditioning

### v1

@@ 1 Introduction
Prediction under constraint lies at the heart of scientific inquiry. Constraints appear in many forms, moment conditions, invariances, sparsity, or limited observability, and pervade the modelling process in both theory and practice. Whether in econometrics, machine learning, or Bayesian analysis, we often possess only partial information about the system under study. How should one operationalize such partial knowledge in a coherent probabilistic framework? This question motivates our study of Bayesian prediction under moment conditioning.
From a sequential perspective, prediction provides the natural operational meaning of inference. The Bayesian predictive distribution defines the law governing future observations given the past. When this sequence is exchangeable, de Finetti’s representation theorem (see 14, 1, 16, 17 for classic results and 6 for modern information-theoretic extensions) shows that prediction is equivalent to averaging over i.i.d. laws drawn from a latent measure. Our goal is to understand how this predictive law behaves when the underlying sequence is subject to constraints, when moment conditions, partial identifications, or structural symmetries restrict the admissible laws.
We study the concentration of the Bayesian predictive distribution toward a single law determined by the constraint geometry. The phenomenon, which we term predictive collapse, captures how partial information causes predictive uncertainty to contract in a way governed by curvature. In this paper, we develop a finite, curvature-adaptive framework that explains this contraction without invoking any large-deviation principle.
@@ 1.1 Connections with Prior Work
Our approach builds on the literature on partial-information models and constrained inference, including empirical likelihood [45, 46], exponentially tilted likelihood [51], and proper Bayesian likelihoods [40]. Following the information-theoretic program of Jaynes [28], we interpret Bayesian updating under constraints as an information projection onto the space of admissible laws. This leads naturally to the Heath–Sudderth finite representation [24], where prediction arises from a discrete prior supported on admissible empirical measures. Classical constructions such as empirical likelihood and Bayesian empirical likelihood correspond to alternative specifications of this prior geometry; our analysis provides the unifying curvature-based principle governing their contraction behavior.
From a probabilistic standpoint, the framework connects the perspectives of de Finetti and Sanov. Doob [18] identified posterior convergence as a martingale property, while Sanov’s theorem [49] establishes concentration of empirical measures via large deviations. Recent work by Polson and Zantedeschi [47] synthesized these perspectives through parameter emergence in exchangeable experiments; we extend their study by studying the dual phenomenon of predictive collapse, deriving finite-sample, curvature-controlled bounds on the contraction of uncertainty.
@@ 1.2 Contributions
On the theoretical side. We derive a finite-sample analogue of local asymptotic normality (LAN) in the sense of Le Cam and Hájek [35], specialized to constrained exchangeable sequences. Conditioning on moment restrictions induces a Gaussian weighting on the tangent space of the constraint manifold, with curvature determined by the information-geometric Hessian $H^{*}$. This construction reproduces, in finite form, the local quadratic expansion underlying the Le Cam–Hájek theory [36], establishing predictive collapse as the finite counterpart of asymptotic contiguity. We derive explicit total-variation bounds showing that curvature, rather than parametrization, governs predictive contraction.
On the methodological side. Our framework provides both new finite-sample theory and a methodological synthesis. It clarifies how classical approaches, empirical likelihood (EL), Bayesian empirical likelihood (BEL), exponentially tilted empirical likelihood (ETEL and BETEL), generalized method of moments (GMM), and generalized estimating equations (GEE), can be understood through their implied prior geometry on admissible empirical measures. We show that BETEL coincides with the conditional likelihood in the LAN limit experiment, GMM is its optimal quadratic expansion, and GEE inherits robustness from effective curvature. By linking behavior to information curvature, the analysis provides a coherent guide for selecting and tuning constrained inferential procedures, unifying large-deviation, empirical-likelihood, and Bayesian perspectives within a single predictive geometry. The work extends to misspecification via Berk’s theorem, where curvature determines both posterior concentration rates and inference brittleness.
@@
On the practical side. The curvature-based representation provides a quantitative diagnostic for identification strength and model robustness. Computing $\lambda_{\min}(H^{*})$ directly indicates whether moment conditions are strongly or weakly identifying: small values signal flat geometry, slow learning, and robustness; large values indicate sharp curvature, rapid concentration, and potential brittleness. The same parameter guides calibration of empirical-likelihood weights, moment penalties in GMM, and tempered posteriors. Sample-size requirements scale as $n\gtrsim m^{2}\log(1/\varepsilon)/(\varepsilon^{2}\lambda_{\min}(H^{*}))$ to achieve predictive precision $\varepsilon$, providing concrete guidance for experimental design. Together, these insights establish predictive collapse as a general law of constrained statistical inference and a practical guide for uncertainty quantification under partial information.
@@ 2 Bayesian Prediction Under Constraints
Let $\mathcal{X}=\{1,\dots,|\mathcal{X}|\}$ denote a finite sample space with $|\mathcal{X}|\geq 2$, and let $Q$ be a reference distribution satisfying $Q(x)>0$ for all $x\in\mathcal{X}$. We observe independent draws
$X_{1},\dots,X_{n}\;\overset{\text{iid}}{\sim}\;Q,\qquad P_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}},$
where $P_{n}$ is the empirical measure. Although discrete, this formulation encompasses numerous applications: categorical surveys, multinomial outcomes in clinical trials, finite partitions in nonparametric Bayes, or discretizations of continuous phenomena. While our development proceeds in a finite, exchangeable setting, approximate and finite de Finetti representations justify this reduction and quantify errors; see Aldous [1], Diaconis and Freedman [16, 17].
@@ 2.1 Constraint Structure and Geometry
Writing $A=[h(1)\ \cdots\ h(k)]\in\mathbb{R}^{d\times k}$, we can express
@@
When $A$ has rank d<k1d<k-1 and $\alpha$ lies in the convex hull of $\{h(x):x\in\mathcal{X}\}$, the feasible set $E$ forms a smooth (k1d)(k-1-d)-dimensional submanifold of the simplex. Most practical constraints, means, variances, correlations, or inequality constraints through slack variables, are affine or reducible to affine form, making this class broad yet analytically tractable.
The information projection of $Q$ onto $E$,
$P^{*}=\operatorname*{arg\,min}_{P\in E}D_{D}(P\|Q),\qquad D_{D}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\frac{P(x)}{Q(x)},$
balances fidelity to $Q$ with adherence to the constraints and serves as the deterministic benchmark for prediction. Its local geometry governs the finite-sample behaviour of constrained posteriors.
Tangent space and projected Hessian. Because probabilities satisfy $\sum_{x}P(x)=1$, the feasible perturbations around $P^{*}$ lie in the tangent space
@@
of dimension r=k1dr=k-1-d. On this subspace, the ambient Hessian of the Kullback–Leibler divergence, $\nabla^{2}D_{D}(P\|Q)=\mathrm{diag}(1/P)$, must be projected onto $\mathsf{T}^{*}$:
@@
where $\Pi_{\mathsf{T}^{*}}$ is the orthogonal projection matrix onto $\mathsf{T}^{*}$ under the Euclidean metric. Equivalently, if $V$ is an orthonormal basis of $\mathsf{T}^{*}$, then
@@
This operator defines the restriction of the Fisher–Rao metric to the constraint manifold [43, 44], thereby acting as the information-geometric curvature tensor of $E$ at $P^{*}$. From a convex-analytic perspective, this projection corresponds to the I-divergence minimization problem of Csiszár [13], while in the differential-geometric view it coincides with the dual affine connection structure of information geometry [3, 2]. These perspectives together justify interpreting $H^{*}$ as the local information-curvature operator governing predictive contraction.
Curvature interpretation. The smallest eigenvalue $\lambda_{\min}(H^{*})$ quantifies constraint curvature. Large $\lambda_{\min}(H^{*})$ implies sharp curvature, strong identification and rapid predictive concentration, whereas small $\lambda_{\min}(H^{*})$ corresponds to flatter geometry, weaker identification, and slower contraction. This curvature constant will govern all finite-sample rates in the sequel, serving as the quantitative link between constraint geometry and predictive behaviour.
@@ 2.2 Predictive Representation
Let $E_{n}=E\cap\mathcal{P}_{n}$ denote the feasible set of empirical types, where $\mathcal{P}_{n}$ is the set of $n$-point empirical distributions. The aim is to predict $m=o(n)$ future observations $X_{n+1},\dots,X_{n+m}$ under the conditioning event that the past data satisfy the constraint. The Bayesian predictive law is therefore
@@
By the Heath–Sudderth representation [24],
@@
where the first factor is hypergeometric sampling without replacement from $P$, and the second is the conditional weight of each feasible empirical type.
The central question is how close this predictive law $\mu_{n,m}$ is to the benchmark $(P^{*})^{\otimes m}$. The analysis to follow shows that the answer depends quantitatively on the curvature parameter $\lambda_{\min}(H^{*})$, which encapsulates the geometric sharpness of the constraint manifold and thus the rate of predictive collapse.
@@ 3 Main Results
Predictive collapse arises because, once structural constraints are imposed, the posterior predictive distribution places nearly all its probability on empirical types close to the information projection $P^{*}$. The rate of this concentration is determined by the curvature of the constraint manifold, quantified by the smallest eigenvalue $\lambda_{\min}(H^{*})$ of the Kullback–Leibler Hessian. Large curvature produces fast concentration; flat geometry slows it.
@@ 3.1 Geometric mechanism
The Heath–Sudderth representation expresses the constrained predictive law as a mixture over feasible empirical types $P\in E_{n}$, weighted by their multinomial likelihood. Expanding the Kullback–Leibler divergence around $P^{*}$ gives
@@
so the weights behave approximately like a Gaussian density on the tangent space $\mathsf{T}^{*}$ with precision $nH^{*}$. Types farther than order $n^{-1/2}$ from $P^{*}$ have exponentially small probability.
3.2 Curvature-Adaptive Localization
To control the discrete–continuous approximation error, we introduce a curvature-adaptive window around $P^{*}$:
$\rho_{n}=\Bigl(\frac{2\log n}{\lambda_{\min}(H^{*})}\Bigr)^{1/2},\qquad\mathcal{W}_{n}=\bigl\{P\in E_{n}:\|P-P^{*}\|_{2}\leq\rho_{n}n^{-1/2}\bigr\}.$
The window is curvature-adaptive: its radius expands when the constraint manifold is flat (small $\lambda_{\min}$) and contracts under strong curvature. This construction follows the localization method of Lanford [33, 34] for bridging discrete combinatorics and continuous approximations, further developed in modern large-deviation treatments such as Ellis [19], here applied to constrained exchangeable sequences. The formal operator-theoretic characterization, which shows $\rho_{n}$ emerges as a fixed point of a regularization operator, is developed in Appendix A.

### v2

@@ 1 Introduction
Prediction under constraint lies at the heart of scientific inquiry. Constraints arise in many forms—moment conditions, invariances, sparsity, or limited observability—and pervade modelling practice across the sciences. Whether in econometrics, machine learning, or Bayesian analysis, we often possess only partial information about the system under study. How should such knowledge be operationalized within a coherent probabilistic framework? This question motivates our study of Bayesian prediction under moment conditioning.
From a sequential viewpoint, prediction provides the operational meaning of inference: the Bayesian predictive distribution describes the law governing future observations given the past. When this sequence is exchangeable, de Finetti’s representation theorem (see 14, 1, 16, 17 for classical results and 6 for modern information–theoretic extensions) shows that prediction amounts to averaging over i.i.d. laws drawn from a latent measure. Our goal is to understand how this predictive law behaves when the underlying sequence is subject to constraints—when moment conditions, partial identification, or structural symmetries restrict the admissible laws.
We study the concentration of the Bayesian predictive distribution toward a single law determined by the geometry of these constraints. We refer to this phenomenon as predictive collapse: partial information causes predictive uncertainty to contract at a rate governed by curvature. The paper develops a finite, curvature–adaptive framework that explains this contraction directly, without appealing to asymptotic or large–deviation limits, thereby extending the logic of Bayesian coherence to the geometry of constrained prediction.
@@ 1.1 Connections with Prior Work
Our approach builds on the literature on partial-information models and constrained inference, including empirical likelihood [49, 50], exponentially tilted likelihood [54], and proper Bayesian likelihoods [45]. Following the information-theoretic program of Jaynes [28], Bayesian updating under constraints is interpreted as an information projection onto the space of admissible laws. This connects to the Heath–Sudderth representation [24], where prediction arises from a discrete prior supported on admissible empirical measures. Classical formulations, empirical likelihood (EL; 49), Bayesian empirical likelihood (BEL; 55), and Bayesian exponentially tilted empirical likelihood (BETEL; 11), correspond to alternative choices of this prior geometry. Our analysis provides the unifying curvature-based principle governing their finite-sample contraction behavior.
From a probabilistic standpoint, the framework extends the tradition inaugurated by Bruno de Finetti. In his foundational papers [14], de Finetti established that exchangeability implies the existence of a directing measure for predictive distributions. Heath and Sudderth [24] formalized this representation in finite form, and subsequent developments by Diaconis, Freedman, and Aldous [16, 17, 1] linked it to Bayesian nonparametrics and empirical-process theory. More recently, Berta et al. [6] derived a finite de Finetti bound via information-theoretic methods, further enriching this line of work. Our construction adapts these ideas to constrained exchangeable sequences, where moment restrictions act as geometric deformations of the directing measure and predictive curvature replaces Fisher information as the local measure of learning strength.
Finally, this geometric view unites the large-deviation and martingale perspectives of Sanov and Doob. Doob [18] interpreted posterior convergence as a martingale property, while Sanov’s theorem [53] characterized empirical-measure concentration through information divergence. Recent work by Polson and Zantedeschi [51] synthesized these strands through parameter emergence in exchangeable experiments; we extend their analysis to the dual phenomenon of predictive collapse, deriving finite–sample, curvature–controlled bounds on the contraction of uncertainty.
@@ 1.2 Contributions
On the theoretical side. We work throughout with a finite sample space $\mathcal{X}=\{1,\dots,k\}$ and identify each probability distribution $P$ on $\mathcal{X}$ with its probability vector $(P(1),\dots,P(k))$ in the simplex $\Delta_{k}$. Given moment constraints $\mathbb{E}_{P}[h(X)]=\alpha$, we write $P^{*}$ for the information projection (KL-minimizer) of a baseline law $Q$ onto the constraint set, $H^{*}$ for the Hessian of the KL divergence restricted to the tangent space of the constraint manifold at $P^{*}$, and $\lambda_{\theta}$ for the Lagrange multiplier enforcing the constraints. Precise definitions appear in Section 2.
We derive a finite-sample analogue of local asymptotic normality (LAN) in the sense of Le Cam and Hájek [39, 41], specialized to constrained exchangeable sequences. Starting from the tilted de Finetti representation,
$\mathcal{L}(X_{1:m}\mid P_{n}\in E_{n})=\int P^{\otimes m}\,\pi_{n}(dP\mid E_{n}),$
Theorem 3.1 (Gaussian de Finetti) shows that conditioning on empirical moment restrictions induces an exponential tilt on the directing measure,
$\pi_{n}(dP\mid E_{n})\;\approx\;\exp\!\big\{-\tfrac{n}{2}(P-P^{*})^{\top}H^{*}(P-P^{*})\big\}\,dP,$
whose local expansion yields a Gaussian weighting on the tangent space of the constraint manifold. Corollary 4.1 (Finite LAN Limit) identifies the curvature of this weighting with the information–geometric Hessian $H^{*}$, providing a discrete predictive analogue of the Fisher information in the LAN experiment. This reproduces, in finite form, the local quadratic expansion underlying the Le Cam–Hájek theory and suggests predictive collapse as the finite counterpart of asymptotic contiguity. Building on these results, Theorem 5.1 (BETEL as Conditional Likelihood) shows that the Bayesian exponentially tilted empirical likelihood arises as the conditional likelihood implied by exchangeability and moment constraints:
$\pi(\theta\mid X_{1:n})\propto\pi(\theta)\,\prod_{i=1}^{n}P_{\theta}^{*}(X_{i}),$
where the exponential tilt $\exp\{\lambda_{\theta}^{\top}h(x)\}$ is absorbed into $P_{\theta}^{*}$. Together, these results demonstrate that curvature—rather than parametrization—governs the rate and stability of predictive contraction, providing the theoretical foundation for all subsequent methodological results.
On the methodological side. Our framework extends these theoretical results to unify and reinterpret classical inferential procedures through their implied curvature geometry. It clarifies how empirical likelihood (EL), Bayesian empirical likelihood (BEL), exponentially tilted empirical likelihood (ETEL and BETEL), generalized method of moments (GMM), and generalized estimating equations (GEE) arise as specializations of the same conditional predictive law. Theorem 5.3 (Quadratic Expansion of the Conditional Likelihood) shows that GMM represents the optimal quadratic approximation to BETEL in the LAN limit,
$\mathcal{L}_{\mathrm{cond}}(\theta)\;\approx\;\exp\!\Big[-\tfrac{n}{2}\,\bar{g}_{n}(\theta)^{\top}\big(A_{\theta}H_{\theta}^{*-1}A_{\theta}^{\top}\big)^{-1}\bar{g}_{n}(\theta)\Big],$
while Theorem 5.4 (Curvature–Robust Estimating Equations) demonstrates that GEE inherits its robustness from effective curvature adjustments within the tangent space. Finally, Theorem 6.1 (Predictive Misspecification) extends the analysis to model misspecification via Berk’s theorem [5], showing that curvature determines both posterior concentration rates and the brittleness of inference under approximate constraints. Together, these results provide a single predictive geometry that unifies large-deviation, empirical-likelihood, and Bayesian approaches.
@@
On the practical side. The curvature–based representation offers a quantitative diagnostic for identification strength and model robustness. Computing $\lambda_{\min}(H^{*})$ directly reveals whether moment conditions are strongly or weakly identifying: small values indicate flat geometry, slow learning, and robustness; large values correspond to sharp curvature, rapid concentration, and potential brittleness. The same parameter guides calibration of empirical-likelihood weights, moment penalties in GMM, and tempered posteriors. Sample–size requirements scale as
$n\gtrsim\frac{m^{2}\log(1/\varepsilon)}{\varepsilon^{2}\lambda_{\min}(H^{*})}$
to achieve predictive precision $\varepsilon$, providing concrete guidance for experimental design. Together, these insights position predictive collapse as a unifying principle for constrained statistical inference and a practical tool for uncertainty quantification under partial information.
Scope and conventions. The results of this paper are established rigorously for finite sample spaces $\mathcal{X}=\{1,\dots,k\}$ under exchangeability and affine moment constraints. Within this setting, Theorems 3.1 and 3.4 provide non-asymptotic bounds with explicit dependence on the curvature parameter $\lambda_{\min}(H^{*})$. The connections to BETEL, GMM, GEE, and Berk’s theorem (Theorems 5.16.1) are derived via local quadratic approximation and hold to leading order in the LAN regime. The broader geometric interpretations—linking predictive collapse to Le Cam theory, information geometry, and large-deviation principles—are offered as a unifying perspective rather than as standalone theorems; they are supported by the formal results but are not claimed to extend beyond the finite, exchangeable setting without additional work.
Figure 1 provides an illustration of the steps taken in the paper to obtain our results.
[…]
Figure 1: Constraint geometry inside the probability simplex $\Delta_{k}$, adapted from the information–geometric visualizations of Kass and Vos [31], Kass [30]. The feasible manifold $E$ represents the distributions satisfying $\mathbb{E}_{P}[h(X)]=\alpha$. The information projection $P^{*}$ is where a Kullback–Leibler contour centered at the baseline model $Q$ is tangent to $E$. The tangent space $\mathsf{T}^{*}$ and the local curvature encoded by the projected Hessian $H^{*}$ govern the Gaussian weighting in Theorem 3.1, quantifying predictive contraction via $\lambda_{\min}(H^{*})$.
@@ 2 Bayesian Prediction Under Constraints
Let $\mathcal{X}=\{1,\dots,k\}$ be a finite sample space with $k\geq 2$ elements. We observe an exchangeable sequence
$X_{1},\dots,X_{n},\qquad P_{n}=\frac{1}{n}\sum_{i=1}^{n}\delta_{X_{i}},$
where $P_{n}$ is the empirical measure. Let $Q$ be a baseline (reference) distribution on $\mathcal{X}$ satisfying $Q(x)>0$ for all $x$. When the data-generating law coincides with $Q$, the setting is well-specified; when it does not, the framework reduces to the misspecified analysis of Section 6, with $Q$ serving as the anchor for the KL projection. Throughout, we identify each probability distribution $P$ on $\mathcal{X}$ with its probability vector $(P(1),\dots,P(k))\in\mathbb{R}^{k}$; all vector operations on $P$ are to be understood in this sense. Although discrete, this formulation encompasses numerous applications: categorical surveys, multinomial outcomes in clinical trials, finite partitions in nonparametric Bayes, or discretizations of continuous phenomena. While our development proceeds in a finite, exchangeable setting, approximate and finite de Finetti representations justify this reduction and quantify errors; see Aldous [1], Diaconis and Freedman [16, 17].
@@ 2.1 Constraint Structure and Geometry
Writing $A=[h(1)\ \cdots\ h(k)]\in\mathbb{R}^{d\times k}$, we may express the feasible set as the intersection
@@
When $A$ has rank d<k1d<k-1 and $\alpha$ lies in the convex hull of $\{h(x):x\in\mathcal{X}\}$, $E$ forms a smooth (k1d)(k-1-d)-dimensional submanifold of the simplex. Most constraints of practical interest—means, variances, correlations, or inequalities handled via slack variables—are affine or reducible to affine form, making this geometry broad yet analytically tractable.
Information projection. Given a baseline model $Q$, the information projection onto $E$ is defined by
$P^{*}=\operatorname*{arg\,min}_{P\in E}D_{D}(P\|Q),\qquad D_{D}(P\|Q)=\sum_{x\in\mathcal{X}}P(x)\log\frac{P(x)}{Q(x)}.$
The point $P^{*}$ balances fidelity to $Q$ with adherence to the constraints and serves as the deterministic benchmark for prediction. Its local curvature will determine the rate and shape of finite-sample contraction.
Tangent space and projected Hessian. Because $\sum_{x}P(x)=1$, feasible perturbations around $P^{*}$ lie in the tangent space
@@
of dimension r=k1dr=k-1-d. On this subspace, the ambient Hessian of the Kullback–Leibler divergence, $\nabla^{2}D_{D}(P\|Q)=\mathrm{diag}(1/P)$, is projected to yield the effective curvature operator
@@
where $\Pi_{\mathsf{T}^{*}}$ denotes the orthogonal projection onto $\mathsf{T}^{*}$. If $V$ is an orthonormal basis for $\mathsf{T}^{*}$, then
@@
This operator represents the Fisher–Rao metric restricted to the constraint manifold [48] and functions as the local information curvature tensor of $E$ at $P^{*}$. From a convex–analytic viewpoint, this corresponds to the I-divergence projection of Csiszár [13]; from the differential–geometric view, it coincides with the dual affine connection of information geometry [3, 2]. Together these perspectives justify interpreting $H^{*}$ as the fundamental local operator governing predictive contraction.
Curvature interpretation. The smallest eigenvalue $\lambda_{\min}(H^{*})$ quantifies constraint curvature. Large $\lambda_{\min}(H^{*})$ implies sharp curvature, strong identification, and rapid predictive concentration, whereas small $\lambda_{\min}(H^{*})$ corresponds to flatter geometry, weaker identification, and slower contraction. This curvature constant will appear throughout the paper as the single geometric parameter linking constraint structure and predictive learning.
@@ 2.2 Predictive Representation
Let $\mathcal{X}$ be a finite observation space, and let $\mathcal{P}$ denote the simplex of all probability measures on $\mathcal{X}$. For each sample size $n$, let $\mathcal{P}_{n}\subset\mathcal{P}$ denote the set of empirical (type) distributions supported on at most $n$ points, and let $E=\{P\in\mathcal{P}:\mathbb{E}_{P}[h(X)]=\alpha\}$ be the population constraint manifold induced by a bounded feature map $h:\mathcal{X}\to\mathbb{R}^{d}$ and target moment $\alpha\in\mathbb{R}^{d}$. The feasible empirical types are then $E_{n}=E\cap\mathcal{P}_{n}$.
We observe $X_{1:n}=(X_{1},\dots,X_{n})$ and denote by $P_{n}$ its empirical distribution. The goal is to predict $m=o(n)$ future observations $X_{n+1},\dots,X_{n+m}$ under the conditioning event that the past sample satisfies the constraint $P_{n}\in E_{n}$. The corresponding Bayesian predictive law is therefore
@@
By the Heath–Sudderth finite representation [24],
@@
where the first factor is the hypergeometric sampling probability of drawing $m$ observations without replacement from the empirical type $P$, and the second factor gives the conditional weight of each feasible empirical distribution.
The central question is how close the constrained predictive law $\mu_{n,m}$ is to the benchmark product measure $(P^{*})^{\otimes m}$, where $P^{*}$ denotes the information projection of the unconstrained law $Q$ onto $E$, defined as
$P^{*}=\arg\min_{P\in E}D_{D}(P\|Q).$
The analysis to follow shows that this proximity is governed by the curvature parameter $\lambda_{\min}(H^{*})$, the smallest eigenvalue of the information–geometric Hessian
$H^{*}=\Pi_{\mathsf{T}^{*}}\,\nabla^{2}D_{D}(P\|Q)\big|_{P=P^{*}}\,\Pi_{\mathsf{T}^{*}},$
the Hessian of the KL divergence projected onto the tangent space $\mathsf{T}^{*}$ of the constraint manifold at $P^{*}$ (see Section 2.1 for the precise definition). This curvature quantifies the local sharpness of the constraint manifold and thus determines the rate of predictive collapse.
@@ 3 Main Results
Predictive collapse arises because, once structural constraints are imposed, the posterior predictive distribution places nearly all its probability on empirical types close to the information projection $P^{*}$. The rate of this concentration is determined by $\lambda_{\min}(H^{*})$: large curvature implies fast concentration, while flat geometry yields slower learning and greater robustness.
@@ 3.1 Geometric Mechanism
The Heath–Sudderth representation expresses the constrained predictive law as a mixture over feasible empirical types $P\in E_{n}$, weighted by their multinomial likelihoods. Expanding the Kullback–Leibler divergence around $P^{*}$ gives
@@
so that, to second order, the weights behave like a Gaussian density on the tangent space $\mathsf{T}^{*}=\{v\in\mathbb{R}^{k}:\mathbf{1}^{\top}v=0,\;Av=0\}$ with precision matrix $nH^{*}$. Empirical types farther than order $n^{-1/2}$ from $P^{*}$ therefore have exponentially small probability.
3.2 Curvature–Adaptive Localization
To control the discrete–continuous approximation error, we introduce a curvature–adaptive localization window around $P^{*}$,
$\rho_{n}=\Bigl(\frac{2\log n}{\lambda_{\min}(H^{*})}\Bigr)^{1/2},\qquad\mathcal{W}_{n}=\bigl\{P\in E_{n}:\|P-P^{*}\|_{2}\leq\rho_{n}/\sqrt{n}\bigr\}.$
The effective radius $\rho_{n}/\sqrt{n}$ expands when the constraint manifold is flat (small $\lambda_{\min}$) and contracts under strong curvature. This construction follows the localization method of Lanford [35, 36] for bridging discrete combinatorics and continuous Gaussian approximations, as further developed in the large–deviation analysis of Ellis [19]. Here it is applied to constrained exchangeable sequences. A formal operator–theoretic characterization, showing that $\rho_{n}$ arises as a fixed point of a regularization operator, is provided in Appendix A.